# Q-Net: Transferable Queue Length Estimation via Kalman-based Neural Networks

### v1

@@ Abstract
Estimating queue lengths at signalized intersections remains a challenge in traffic management, especially under partially observed conditions where vehicle flows are not fully captured. This paper introduces Q-Net, a data-efficient and interpretable framework for queue length estimation that performs robustly even when traffic conservation assumptions are violated. Q-Net integrates two widely available and privacy-friendly data sources: (i) vehicle counts from loop detectors near stop lines, and (ii) aggregated floating car data (aFCD), which divides each road section into segments and provides segment-wise average speed measurements. These data sources often differ in spatial and temporal resolution, creating fusion challenges. Q-Net addresses this by employing a tailored state-space model and an AI-augmented Kalman filter, KalmanNet, which learns the Kalman gain from data without requiring prior knowledge of noise covariances or full system dynamics. We build on the vanilla KalmanNet pipeline to decouple measurement dimensionality from section length, enabling spatial transferability across road segments. Unlike black-box models, Q-Net maintains physical interpretability, with internal variables linked to real-world traffic dynamics. Evaluations on main roads in Rotterdam, the Netherlands, demonstrate that Q-Net outperforms baseline methods by over 60% in Root Mean Square Error (RMSE), accurately tracking queue formation and dissipation while correcting aFCD-induced delays. Q-Net also demonstrates strong spatial and temporal transferability, enabling deployment without costly sensing infrastructure like cameras or radar. Additionally, we propose a real-time variant of Q-Net, highlighting its potential for integration into dynamic, queue-based traffic control systems.
@@ 1 Introduction
Traffic queues raise operational, safety, and environmental concerns that necessitate effective monitoring and control systems. Congestion in cities reduces network throughput and limits agglomeration economies (Sweet, 2011). Beyond efficiency concerns, long queues increase accident likelihood (Pasquale et al., 2018), elevate noise levels (Chevallier et al., 2009), contribute to driver stress and mental health issues (Nadrian et al., 2019), and generate localized air pollution due to stationary vehicles (Matzoros and Van Vliet, 1992). Real-time queue length estimation addresses these challenges by enabling adaptive traffic control strategies that dynamically optimize signal timing in response to changing conditions, while accurate queue length estimation more broadly supports the development of advanced queue-based control algorithms and network-wide optimization strategies (Li et al., 2021; Hoogendoorn et al., 2016). However, despite being a longstanding focus in traffic management, queue length estimation remains challenging due to incomplete sensor coverage, unpredictable driver behavior, and intricate network topologies.
@@
Queue length estimation methods at signalized intersections are characterized by their type of inputs. These input data can be Lagrangian and Eulerian types. Lagrangian data are collected from mobile sensors that move with traffic flow, primarily through GPS-equipped probe vehicles. Recent research has extensively explored probe vehicle trajectory data for queue length estimation, employing methodologies ranging from shockwave theory to probabilistic likelihood-based approaches (Cheng et al., 2011; Hu et al., 2022; Zhao et al., 2019; Cheng et al., 2012; Zhao et al., 2021). However, these methods typically require prior knowledge of probe vehicle penetration rates and raise privacy concerns, as participants must reveal their positions to external traffic monitoring systems (Rose, 2006).
In contrast, Eulerian data capture traffic conditions over fixed spatial segments, coming from either stationary sensors or aggregated mobile sources. Video-based systems, enhanced by computer vision, can detect queue ends (Umair et al., 2021; Albiol et al., 2011), though their performance may degrade under poor lighting or bad weather. Radar systems, while capable of providing direct queue measurements, are expensive to install and maintain (Bernas et al., 2018). Loop detectors are among the most widely used sensors due to their ability to provide traffic counts, occupancy, and speed measurements at a low cost (MÜCK, 2002; Liu et al., 2009; Geroliminis and Skabardonis, 2011; Abewickrema et al., 2023). Recently, aggregated floating car data (aFCD) from commercial vendors like INRIX, HERE, Be-Mobile, and TomTom represents another Eulerian source. These data are anonymized and aggregated over road segments and time intervals to comply with privacy regulations, offering advantages in terms of data availability without requiring high penetration rates (van der Loop et al., 2017). The aFCD is of widespread use in France (Gitahi et al., 2020), Turkey (Altıntaşı, 2018), and the Netherlands (Mobiliteitsscan, ).
Despite advancements in high-resolution video analytics and individual-level probe vehicle tracking, many cities still rely on conventional data sources, such as loop detectors, due to infrastructure constraints, limited budgets, and privacy regulations. Given these practical considerations, this study focuses on a queue length estimation that uses two widely available data sources as input: loop detector measurements and aFCD, which ensures broad applicability across diverse urban settings and aligns well with infrastructure where these data are readily available.
@@ 1.1 Literature review
Queue length estimation at signalized intersections can generally be categorized into four approaches: the input-output models (Webster, 1958; Sharma et al., 2007), shockwave theory (Lighthill and Whitham, 1955; Liu et al., 2009), speed-based methods (Dinh et al., 2014), and filtering algorithms (Yin et al., 2018).
Input-output models estimate queue length based on the conservation equation of cumulative traffic inflow and outflow at signalized sections. Traditionally, researchers have relied on vehicle counts recorded by road detectors to characterize the queuing process (Webster, 1958). Later, Vigos et al. (2008) enhanced this approach by applying real-time flow and occupancy measurements to adjust queue length estimates in accordance with conservation laws. While the input-output model is simple and straightforward, it does not capture the spatiotemporal dynamics of queues and is highly sensitive to the data measurement quality. More recent studies have improved input-output models by integrating probe vehicle data, enabling real-time estimation of the queue tail position (Wang et al., 2021; Tang et al., 2023), but these models require a minimal penetration rate.
Shockwave theory describes the queue formation and dissipation dynamics at signalized intersections, with shockwave speed indicating traffic state transitions. Liu et al. (2009) used loop detector data for cycle-wise queue estimation, later refined by Abewickrema et al. (2023) to handle both oversaturated and undersaturated conditions. These methods rely on identifying occupancy breakpoints across signal phases, but in practice, such patterns are often obscured by noise and traffic heterogeneity. To address this, studies such as Cai et al. (2014); Wang et al. (2017), fuse probe vehicle data with detector measurements to estimate shockwave speeds, enabling the estimation of cycle-wise maximum and residual queues. An et al. (2018) proposed a hybrid approach: applying input-output models when breakpoints are unreliable, and using shockwave analysis otherwise. While these shockwave-based models offer a theoretical basis for queue evolution, they assume uniform traffic conditions, which rarely hold in urban settings.
@@
Kalman filtering algorithms are based on state-space models, which describe system dynamics through a hidden state evolution equation and a measurement (observation) equation. The hidden state represents queue-length-related variables that evolve over time but are not directly observable. As is standard in state-space modeling, different data sources can be integrated into either the evolution model or the measurement model, providing flexibility. This framework is crucial for queue estimation because it incorporates noisy, indirect measurements from various sensors while explicitly modeling uncertainties in traffic dynamics and observations. The Kalman filter recursively estimates the hidden states by alternating between prediction—forecasting the system state and expected measurements—and update steps that refine estimates based on actual observations, providing a robust, real-time solution for tracking queue dynamics.
Several studies have adapted Kalman-based methods to queue length estimation. Abewickrema et al. (2023) applied an extended Kalman filter, modeling the cycle-wise maximum and residual queue lengths as the system state, which is assumed to remain static within each signal cycle. The measurement model is derived from shockwave theory to capture phase breakpoints. Hu et al. (2022) implemented two independent extended Kalman filters to separately track queue formation and discharge waves, using a measurement model that estimates shockwave location and speed from the hidden state. Wang et al. (2021) used a standard Kalman filter where the hidden state represents the number of queued vehicles. The evolution model is grounded in input-output theory, while the measurement model estimates queue length by accounting for the probe vehicle penetration rate. Similarly, Heshami and Kattan (2021) employed an evolution model based on input-output theory, but with a measurement model derived from a case-based reasoning algorithm.
Although traditional Kalman filters are interpretable and computationally efficient, they depend on full prior knowledge of system evolution dynamics and measurement models, which often do not match practical scenarios. For example, Noise covariances can be complex, unknown, and time-varying, which can degrade the performance of linear Kalman-based methods (Li et al., 2015). These limitations are particularly pronounced in queue length estimation, where traffic dynamics are inherently nonlinear, influenced by heterogeneous and time-varying factors (Ahmed et al., 2021; Chauhan et al., 2021) such as turning movements, and challenging to model accurately in partially observed traffic scenarios.
@@ 1.2 Research gap and contribution
There are three key limitations in current approaches. First, limitations in problem setup: most existing studies are developed and validated in well-controlled environments, often overlooking practical challenges such as partially observed traffic flows. In real-world urban settings, it is common for vehicles to enter or exit the study area from adjacent roads or unsignalized intersections without being monitored due to sensor installation constraints and failures (Federal Highway Administration, 2006; Shang et al., 2024). Second, data dependency: studies using loop detectors tend to rely on established models like input-output and shockwave theories, which impose strict requirements on data quality (e.g., accurate inflow/outflow counts and clear occupancy patterns) that are frequently unmet in real-world settings. On the other hand, aFCD—valued for its low cost, wide coverage, and privacy advantages—remains underutilized. As a result, current methods lack generalizability across different network sensor setups, with performance degrading significantly when applied to imperfect real world. Third, methodological constraints: many approaches depend on traditional Kalman filters and Extended Kalman filters, which impose linear or linearized models and known noise covariances. These assumptions are often unrealistic for dynamic traffic conditions at intersections, limiting the adaptability and scalability of such methods in diverse real-world applications.
To address these limitations, we propose Q-Net, a novel queue length estimation approach tailored to urban environments. First, Q-Net handles partially observed traffic flows, making it robust to incomplete monitoring coverage in real-world settings. Second, our approach leverages aggregated loop detector and aFCD data as inputs, both of which are widely available and privacy-friendly data sources. Third, to overcome the methodological constraints of traditional Kalman filters while preserving interpretability, we adopt an AI-augmented Kalman filtering approach that combines deep learning with state-space modeling. To achieve this, we adapt KalmanNet (Revach et al., 2022) to the queue length estimation problem. KalmanNet and its variants have gained attention in signal processing for their ability to learn complex dynamics while maintaining the recursive structure and transparency of classical filtering methods (Buchnik et al., 2023; Revach et al., 2023; Choi et al., 2023).
Our main contributions are as follows:
@@
    Interpretable state-space formulation: We develop a state-space model for queue length estimation that relies solely on loop detector data installed near the stop lines of adjacent intersections and aggregated Floating Car Data (aFCD) as inputs, both of which are low cost and widely available. While loop detectors can be flexibly positioned along the roadway, our method remains applicable by accounting for the time lag between vehicle detection and arrival at the stop line. The state-space model consists of an evolution model based on cumulative vehicle counts from loop detector data and a measurement model derived from aFCD speeds. It accommodates general traffic conditions including scenarios with partially observed traffic flows. This design enables effective fusion of data sources with differing spatial and temporal resolutions.
@@
    Data-driven solution: Q-Net addresses the trade-off between modeling capacity and interpretability of black-box neural networks. It integrates Kalman filtering principles with deep learning, following AI-augmented Kalman filtering techniques. Specifically, building on the tailored state-space formulation, Q-Net employs neural networks to calculate the Kalman gain and maintains physically meaningful internal states. Q-Net uses a measurement grouping strategy that makes model dimensions independent of road section length, which ensures transferability across different traffic scenarios—the model can be trained on one section and tested on another. Q-Net is highly data-efficient with only 952 parameters and trainable on 13-day datasets. A real-time learning mechanism is also incorporated for adaptive traffic control. To our knowledge, Q-Net represents the first application of AI-augmented Kalman filtering to queue length estimation.
@@
    Real-world validation: Q-Net has been validated on main urban road sections in Rotterdam, the Netherlands. Training on one road section and testing across different days yields over 60% RMSE improvement compared to baseline models, showing strong temporal transferability. The same trained model works directly on adjacent road sections without any finetuning. These adjacent sections present different traffic patterns due to their unique network configurations, yet Q-Net maintains reliable performance validated through quantitative evaluation and visual inspection.
The remainder of this paper is organized as follows: Section 2 introduces the fundamentals of the Kalman filter. Section 3 presents the problem setup and highlights key challenges. Section 4 describes how Q-Net models traffic dynamics using a state-space formulation, and Section 5 details its filtering procedure. Section 6 presents the implementation of Q-Net on major road sections in Rotterdam, analyzes the results, demonstrates its transferability, and showcases its real-time version for traffic control. Finally, Section 7 concludes the paper.
@@ 2 Background in Kalman filtering
State-space models describe how a system’s internal (latent) state and measurements evolve across time steps. Let $\mathbf{x}_{t}$ denote the latent state of the system at time $t$, $\mathbf{u}_{t}$ the control input, and $\mathbf{y}_{t}$ the observed measurement. The system is governed by:
(1) State evolution model that describes how the state evolves over time:
@@
(2) Measurement model that relates the observed measurement to the current state:
@@
where $\mathbf{H}_{t}$ is the observation matrix and $\mathbf{v}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{t})$ is zero-mean Gaussian measurement noise. The noise $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ are modeled to be mutually independent and also independent of both the states and the measurements. The initial state $\mathbf{x}_{0}$ is modeled to be Gaussian, with known mean $\hat{\mathbf{x}}_{0}$ and covariance $\hat{\boldsymbol{\Sigma}}_{0}$.
@@
$\boldsymbol{\Sigma}_{t|t-1}$ $=\mathbf{F}_{t}\boldsymbol{\Sigma}_{t-1|t-1}\mathbf{F}_{t}^{\top}+\mathbf{Q}_{t}.$ (4)
Throughout this paper, $(\cdot)^{\top}$ denotes the matrix transpose operation. The predicted measurement is:
@@
$\mathbf{\Sigma}_{t|t}$ $=\mathbf{\Sigma}_{t|t-1}-\mathbf{\mathcal{K}}_{t}\mathbf{S}_{t|t-1}\mathbf{\mathcal{K}}_{t}^{T}.$ (8)
The innovation is given by $\Delta\mathbf{y}_{t}=\mathbf{y}_{t}-\hat{\mathbf{y}}_{t|t-1}$, representing the prediction error between the observed measurement $y_{t}$ and the one-step-ahead prediction $\hat{y}_{t|t-1}$. The Kalman gain $\mathbf{\mathcal{K}}_{t}$ is defined as $\mathbf{\mathcal{K}}_{t}=\mathbf{\Sigma}_{t|t-1}\mathbf{H}^{T}\mathbf{S}_{t|t-1}^{-1}$. It is chosen to minimize the mean squared error (MSE) of the state estimate. The classical Kalman filter thus takes the posteriori first- and second-order moments from the previous time step and produces those for the current step, enabling online, recursive state estimation. Another widely used extension is the Extended Kalman Filter (EKF), which handles nonlinear systems by locally linearizing the state and measurement models via their Jacobian matrices.

### v2

# Q-Net: Queue Length Estimation via Kalman-based Neural Networks

@@ Abstract
Estimating queue lengths at signalized intersections is a long-standing challenge in traffic management. Partial observability of vehicle flows complicates this task despite the availability of two privacy preserving data sources: (i) aggregated vehicle counts from loop detectors near stop lines, and (ii) aggregated floating car data (aFCD) that provide segment-wise average speed measurements. However, how to integrate these sources with differing spatial and temporal resolutions for queue length estimation is rather unclear. Addressing this question, we present Q Net: a robust queue estimation framework built upon a state-space formulation. This formulation addresses key challenges in queue modeling, such as violations of traffic conservation assumptions. To overcome the limitations of standard filtering models in integrating diverse data sources, Q-Net employs an AI-augmented Kalman filter for estimation. Q-Net follows the Kalman predict-update framework and maintains physical interpretability, with internal variables linked to real-world traffic dynamics. Q-Net can be implemented in real-time, making it suitable for integration into queue-based traffic control systems. To achieve spatial transferability across road sections, we group aFCD measurements into fixed-size groups. This strategy ensures the dimension of Q-Net’s learnable parameters is independent of section length. Evaluations on urban main roads in Rotterdam, the Netherlands, show that Q-Net outperforms baseline methods, accurately tracking queue formation and dissipation while correcting aFCD-induced delays. By combining data efficiency, interpretability, and strong transferability, Q Net makes accurate queue length estimation possible without costly sensing infrastructure like cameras or radar.
@@ 1 Introduction
Traffic queues raise operational, safety, and environmental concerns that necessitate effective monitoring and control systems. Congestion in cities reduces network throughput and limits agglomeration economies (Sweet, 2011). Long queues increase also accident likelihood (Pasquale et al., 2018), elevate noise levels (Chevallier et al., 2009), contribute to driver stress and mental health issues (Nadrian et al., 2019), and generate localized air pollution (Matzoros and Van Vliet, 1992). Real-time queue length estimation addresses these challenges by enabling traffic control strategies that adapt signal timing in response to changing conditions, while accurate queue length estimation more broadly supports the development of advanced queue-based control algorithms and network-wide optimization strategies (Li et al., 2021; Hoogendoorn et al., 2016). However, queue length estimation remains challenging due to incomplete sensor coverage, limited data availability, unpredictable driver behavior, and intricate network topologies.
@@
Queue length estimation methods at signalized intersections are characterized by their type of inputs such as Lagrangian and Eulerian types. Lagrangian data are collected from mobile sensors that move with traffic flow, primarily through GPS-equipped probe vehicles. Recent research has explored probe vehicle trajectory data for queue length estimation, employing methodologies ranging from shockwave theory to probabilistic likelihood-based approaches (Cheng et al., 2011; Hu et al., 2022; Zhao et al., 2019; Cheng et al., 2012; Zhao et al., 2021). These methods typically require prior knowledge of probe vehicle penetration rates and raise privacy concerns, as participants must reveal their positions to external monitoring systems (Rose, 2006). In contrast, Eulerian data capture traffic conditions over fixed spatial segments, coming from either stationary sensors or aggregated mobile sources. Computer vision-based systems can detect queue ends (Umair et al., 2021; Albiol et al., 2011), though their performance may degrade under poor lighting or bad weather. Radar systems, while capable of providing direct queue measurements, are expensive to install and maintain (Bernas et al., 2018).
Loop detectors remain among the most widely used traffic sensors because of their high accuracy in measuring vehicle counts, occupancy, and speed. Although installation requires road surface disruption, their advantage lies in the extensive existing infrastructure (MÜCK, 2002; Liu et al., 2009; Geroliminis and Skabardonis, 2011; Abewickrema et al., 2023). Many cities continue to rely on these conventional Eulerian data sources owing to infrastructure constraints, limited budgets, and strict privacy regulations. Recently, aggregated floating car data (aFCD) from commercial vendors like INRIX, HERE, Be-Mobile, and TomTom represents another Eulerian source. These data are anonymized and aggregated over road segments and time intervals to comply with privacy regulations, offering advantages in data availability without requiring high penetration rates (van der Loop et al., 2017). The aFCD is widespread in France (Gitahi et al., 2020), Turkey (Altıntaşı, 2018), and the Netherlands (Mobiliteitsscan, ). Given these practical considerations, this study focuses on a queue length estimation that uses two complementary data sources as input: loop detector measurements and aFCD, which ensure broad applicability across diverse urban settings and aligns well with infrastructure where these data are readily available. The challenge remains into integrating these diverse type of data in a unified and principled way for queue length estimation.
@@ 1.1 Literature review
Queue length estimation at signalized intersections can be categorized into four approaches: the input-output models (Webster, 1958; Sharma et al., 2007), shockwave theory (Lighthill and Whitham, 1955; Liu et al., 2009), speed-based methods (Dinh et al., 2014), and filtering algorithms (Yin et al., 2018).
Input-output models estimate queue length based on the conservation equation of cumulative traffic inflow and outflow at signalized sections. Traditionally, researchers relied on vehicle counts recorded by road detectors to characterize the queuing process (Webster, 1958). Later, Vigos et al. (2008) enhanced this approach by applying real-time flow and occupancy measurements to adjust queue length estimates. While the input-output model is straightforward, it does not capture the spatiotemporal dynamics of queues and is highly sensitive to the data measurement quality. Recent studies have improved input-output models by integrating probe vehicle data, enabling real-time estimation of the queue tail position (Wang et al., 2021; Tang et al., 2023), but these models require a minimal penetration rate.
Shockwave theory describes the queue formation and dissipation dynamics at signalized intersections, with shockwave speed indicating traffic state transitions. Liu et al. (2009) used loop detector data for cycle-wise queue estimation, later refined by Abewickrema et al. (2023) to handle both oversaturated and undersaturated conditions. These methods identify occupancy breakpoints across signal phases, but such patterns are often obscured by noise and traffic heterogeneity. To address this, studies such as Cai et al. (2014); Wang et al. (2017), fuse probe vehicle data with detector measurements to estimate shockwave speeds, enabling the estimation of cycle-wise maximum and residual queues. An et al. (2018) proposed a hybrid approach: applying input-output models when breakpoints are unreliable, and using shockwave analysis otherwise. While these shockwave-based models offer a theoretical basis for queue evolution, they assume uniform traffic conditions, which rarely hold in urban settings.
@@
Kalman filtering algorithms are based on state-space models that describe the system dynamics through a hidden state evolution equation and a measurement (observation) equation. The hidden state represents queue-length-related variables that evolve over time but are not directly observable. Different data sources can be integrated into either the evolution model or the measurement model, providing flexibility. This framework is crucial for queue estimation because it incorporates noisy, indirect measurements from various sensors while explicitly modeling uncertainties in traffic dynamics and observations. The Kalman filter recursively estimates the hidden states by alternating between prediction—forecasting the system state and expected measurements—and update steps that refine estimates based on actual observations, providing a robust, real-time solution for tracking queue dynamics.
Abewickrema et al. (2023) applied an extended Kalman filter, modeling the cycle-wise maximum and residual queue lengths as the system state, which is assumed to remain static within each signal cycle. The measurement model is derived from shockwave theory to capture phase breakpoints. Hu et al. (2022) implemented two independent extended Kalman filters to separately track queue formation and discharge waves, using a measurement model that estimates shockwave location and speed from the hidden state. Wang et al. (2021) used a standard Kalman filter where the hidden state represents the number of queued vehicles. The evolution model is grounded in input-output theory, while the measurement model estimates queue length by accounting for the probe vehicle penetration rate. Similarly, Heshami and Kattan (2021) employed an evolution model based on input-output theory, but with a measurement model derived from a case-based reasoning algorithm. However, traditional Kalman filters impose linear or linearized system models and depend on full prior knowledge of system evolution dynamics and measurement models, which often do not match practical scenarios. For example, noise covariances can be complex, unknown, and time-varying, which can degrade the performance of linear Kalman-based methods (Li et al., 2015). These limitations are particularly pronounced in queue length estimation, where traffic dynamics are inherently nonlinear, influenced by heterogeneous and time-varying factors (Ahmed et al., 2021; Chauhan et al., 2021) and challenging to model accurately in partially observed traffic scenarios.
@@ 1.2 Research gap and contribution
There are three key limitations in current approaches related to problem setup, data dependency, and methodological constraints. First, related to the limitations in problem setup: most existing studies are developed and validated in well-controlled environments, often overlooking practical challenges such as partially observed traffic flows. In urban settings, vehicles can enter or exit the study area from adjacent roads or unsignalized intersections without being monitored due to sensor installation constraints and failures (Federal Highway Administration, 2006; Shang et al., 2024). Second, related to the data dependency: studies using loop detectors rely on established models like input-output and shockwave theories, which impose strict requirements on data quality (e.g., accurate inflow/outflow counts and clear occupancy patterns) that are frequently unmet in practice. On the other hand, aFCD—valued for its low cost, wide coverage, and privacy advantages—remains underutilized. As a result, current methods lack generalizability across different network sensor setups, with performance degrading significantly when applied to imperfect real world. Third, related to the methodological constraints: many approaches depend on traditional or extended Kalman filters, which impose linearized models and known noise covariances. These assumptions are often unrealistic for dynamic traffic conditions at intersections, limiting their adaptability and scalability in diverse real-world applications.
To address these limitations, we propose Q-Net, a novel queue length estimation approach tailored to urban environments. Q-Net is built upon a state space modeling approach integrated with an AI augmented Kalman filtering framework. First, Q-Net handles partially observed traffic flows, making it robust to incomplete monitoring coverage in real-world settings. Second, Q-Net leverages aggregated loop detector and aFCD data as inputs, both of which are widely available and privacy-friendly data sources. Third, to overcome the methodological constraints of traditional Kalman filters while preserving interpretability, Q-Net exploits deep neural networks to learn the Kalman gain dynamically by modeling time varying and nonlinear noise characteristics. The learning process builds on the KalmanNet principle (Revach et al., 2022), which preserves system dynamics to a large extent and its variants have gained attention in signal processing (Buchnik et al., 2023; Revach et al., 2023; Choi et al., 2023).
Our main contributions are threefold:
@@
    Interpretable state-space formulation: We develop a state-space model for queue length estimation that relies solely on loop detector data installed near the stop lines of adjacent intersections and aFCD as inputs, both of which are low cost and widely available. The state-space model consists of an evolution model based on cumulative vehicle counts from loop detector data and a measurement model derived from aFCD speeds. It accommodates general traffic conditions including scenarios with partially observed traffic flows. This design enables effective fusion of data sources with differing spatial and temporal resolutions.
@@
    Data-driven solution: Q-Net addresses the trade-off between modeling capacity and interpretability in black-box neural networks by integrating Kalman filtering with deep learning, following AI-augmented Kalman filtering principles. Specifically, we develop a problem-specific state space formulation in which the queue length is a latent state, queue change from loop detectors as control input, and aFCD speeds as measurements. Neural networks compute the Kalman gain while maintaining physically meaningful internal states. We employ a measurement grouping strategy considering that measurement covariances relate only to neighboring aFCD speeds. This strategy makes the model dimension independent of road section length, allowing transferability across traffic scenarios, i.e., Q Net can be trained on one section and tested on another.
@@
    Real-world validation: Q-Net has been validated on main urban road sections in Rotterdam, the Netherlands. With only 952 trainable parameters, Q-Net is highly data-efficient and can be trained on a 12-day dataset. It yields improvements over all metrics compared to baseline models, such as Tree-RPROP (Luo et al., 2022) that uses more data sources than Q-Net. The same trained model works directly on adjacent road sections without any finetuning. Despite these adjacent sections presenting different traffic patterns due to their unique infrastructure configurations, Q-Net maintains reliable performance (validated through quantitative evaluation and visual inspection), demonstrating strong spatial transferability. Similarly, applying Q-Net in real-time also achieves promising results for supporting adaptive traffic control. To our knowledge, Q-Net represents the first application of AI-augmented Kalman filtering to queue length estimation.
The remainder of this paper is organized as follows. Section 2 introduces the fundamentals of the Kalman filter. Section 3 presents the problem setup and highlights key challenges. Section 4 describes how Q-Net models traffic dynamics using a state-space formulation, and Section 5 details its filtering procedure. Section 6 presents the implementation of Q-Net on major road sections in Rotterdam, analyzes the results, demonstrates its transferability, and showcases its real-time version for traffic control. Finally, Section 8 concludes the paper.
@@ 2 Background in Kalman filtering
State-space models describe how a system’s internal (latent) state and measurements evolve across time steps. Let $\mathbf{x}_{t}$ denote the latent state of the system at time $t$, $\mathbf{u}_{t}$ the control input, and $\mathbf{y}_{t}$ the observed measurement. The system is governed by a state evolution and a measurement model.
(1) State evolution describes how the state $\mathbf{x}_{t}$ evolves over time:
@@
(2) Measurement model relates the observed measurement to the current state:
@@
where $\mathbf{H}_{t}$ is the observation matrix and $\mathbf{v}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{t})$ is zero-mean Gaussian measurement noise. The noise $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ are mutually independent and also independent of both the states and the measurements. The initial state $\mathbf{x}_{0}$ is as Gaussian with known mean $\hat{\mathbf{x}}_{0}$ and covariance matrix $\hat{\boldsymbol{\Sigma}}_{0}$.
@@
$\boldsymbol{\Sigma}_{t|t-1}$ $=\mathbf{F}_{t}\boldsymbol{\Sigma}_{t-1|t-1}\mathbf{F}_{t}^{\top}+\mathbf{Q}_{t},$ (4)
where $(\cdot)^{\top}$ denotes the matrix transpose operation. The predicted measurement is:
@@
$\mathbf{\mathcal{K}}_{t}$ $=\mathbf{\Sigma}_{t|t-1}\mathbf{H}_{t}^{\top}\mathbf{S}_{t|t-1}^{-1},$ (9)
$\mathbf{\Sigma}_{t|t}$ $=\mathbf{\Sigma}_{t|t-1}-\mathbf{\mathcal{K}}_{t}\mathbf{S}_{t|t-1}\mathbf{\mathcal{K}}_{t}^{\top}.$ (10)
The innovation $\Delta\mathbf{y}_{t}=\mathbf{y}_{t}-\hat{\mathbf{y}}_{t|t-1}$ represents the prediction error between the observed measurement $\mathbf{y}_{t}$ and the one-step-ahead prediction $\hat{\mathbf{y}}_{t|t-1}$. The Kalman gain $\mathbf{\mathcal{K}}_{t}=\mathbf{\Sigma}_{t|t-1}\mathbf{H}_{t}^{\top}\mathbf{S}_{t|t-1}^{-1}$ minimizes the mean squared error (MSE) of the state estimate. It is computed analytically under the assumptions of a linear state-space model with fixed Gaussian noise. The classical Kalman filter thus takes the posteriori first- and second-order moments from the previous time step and produces those for the current step, enabling online, recursive state estimation. Another widely used extension is the Extended Kalman Filter (EKF), which handles nonlinear systems by locally linearizing the state and measurement models via their Jacobian matrices (Maybeck, 1982).